OpenStudy (anonymous):

Please help me with these questions 1. simplify the trigonometric expression. sec theta cos theta A. tan theta B. 1 C. cot theta D. sin theta 2. use the unit circle to find the inverse function value in degrees. tan^-1 square root(3) A. 120 B. 90 C.60 D. 30 3. what values for theta(0

OpenStudy (solomonzelman):

I am going to give you some rules, to HELP you with these questions. For number 1, \(\LARGE\color{blue}{ \bf \frac{1}{cos(x)}=sec(x) }\) and \(\LARGE\color{blue}{ \bf \frac{1}{sec(x)}=cos(x) }\)

OpenStudy (anonymous):

okay and how would you do number 1 like im confused :P

OpenStudy (anonymous):

@rockstargirl17 What answer do u get in 1?

OpenStudy (solomonzelman):

for number 1 re-write the sec(theta) in terms of cos(theta)

OpenStudy (anonymous):

oh okay then after you did that what else do you do ?

OpenStudy (solomonzelman):

At first , after you do that, show me, what are you getting ?

OpenStudy (anonymous):

you would get cos( x) right ?